OpenStudy (anonymous):

In order to take pictures of a dance troupe as it performs, a camera crew rides in a cart on tracks that trace a circular arc. The radius of the arc is 75 ft, and from end to end the cart sweeps out an angle of 172.5 degrees in 20 seconds. Find A.) the length of the track in feet and inches

OpenStudy (anonymous):

the length is about 9.3ft

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

The answer should equal 225.8 ft?

OpenStudy (anonymous):

Make 172.5 degrees into the unit radians. The formula to do so is: \[x \times \frac{ \pi }{ 180 }\] <-- Insert 172.5 in place of x. Formula to get your answer is: arc length = radius*degree in radians Insert what you know into that formula and voila! You get your answer.